The perimeter of a rectangle is 52 inches. if the length of the rectangle is 2 inches more than its width, what are the dimensions of the rectangle.
Answer by erbirdsong(1)

L+L+W+W=2L+2W=52
2(2W)+2W=52
4W+4+2W=52
6W=48
W=8

this above is wrong!
=======================================================================

correct below:

let width = x
so that length is x+2

perimeter is, as said above all 4 sides added:

x + x + x+2 + x+2 = 52
4x + 4 = 52
4x = 48
x = 12

so width = 12
and length = 14
jon.